OpenStudy (anonymous):

What are the solutions of the inequality? 2(b - 8) > 12 Select one: a. b > 6 b. b > 14 c. b > 20 d. b < 20 need help on math homework

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Look : if we have : 2(5+9)=2*5+2*9 Right ? So if have : 2(a+3)=3a+6 we can say : 2(b - 8)=2*b+2*(-8)=2b-16 right ?;) @Bxbba

OpenStudy (anonymous):

b>14

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

@quint41: thank you but I think step by step is better !;)